OpenStudy (anonymous):

can someone help me solve this either using substitution or elimination please? 1/2x + 2/3y = 6 x + y = 10

OpenStudy (anonymous):

You can change the first equation by multiplying it by 6 and this will get rid of the fractions.... (6)1/2x + (6)2/3y = (6)6 3x + 4y = 36 (thats better) x + y = 10 --> x = 10 - y now sub 10 - y in for x in the " new equation ". 3x + 4y = 36 3(10 - y) + 4y = 36 Can you finish it from here ?

OpenStudy (anonymous):

Okay , I was making sure I was on the right track, I managed to eliminate the fractions. What I am having trouble with is the second equation

OpenStudy (whpalmer4):

Which equation is "the second equation" in your eyes, and what is the trouble you are having? Can you show us your work so far?

OpenStudy (anonymous):

the x+y=10

OpenStudy (anonymous):

what i have is 6*1/2x+6*2/3y=6 3x+4y=36

OpenStudy (anonymous):

and I get that x=4, and y=6?

OpenStudy (whpalmer4):

Well, let's check those out: they have to satisfy both equations to be a valid solution to the system: \[x+y=10\] \[4+6=10\checkmark\] \[\frac{1}{2}x+\frac{2}{3}y =6\]\[\frac{1}2(4)+\frac{2}3(6) = 6\]\[2+4=6\checkmark\]Good work!

OpenStudy (anonymous):

o wow, so I had it correct? lol

OpenStudy (whpalmer4):

It's important to check both equations: if you'd made an error somewhere along the line and gotten x=3, y = 7, that would satisfy one equation, but not the other.

OpenStudy (whpalmer4):

Do you want to try doing it by elimination now?

OpenStudy (anonymous):

I can try